# Temperature-Dependent Calibration Procedures for the Silicon Photomultiplier Readout of the Cosmic Ray Veto Detector for the Mu2e Experiment

### v1

@@ 1 Introduction
The Mu2e experiment intends to search for the conversion of a negatively charged muon into an electron in the presence of an aluminum nucleus: [1]. In the Standard Model, this process is forbidden. In the extended Standard Model with non-zero neutrino masses, it proceeds at the experimentally unobservable rate of [2]. Hence, any observation of this process would be unambiguous evidence of new physics.
@@ 2 The Cosmic Ray Veto Detector Design
The CRV scintillator bars are grouped in modules, each module consisting of 64 bars, 16 per layer, with aluminum sheets of 3/8” (9.525 mm) thickness placed between each scintillator layer. The entire CRV consists of 83 total modules. Each scintillator bar is mm2 in cross section, with lengths ranging from 0.985 m to 6.900 m.111The bars have a polystyrene base doped with 1% PPO, and 0.03% POPOP. The bars have a thin TiO2 reflective coating and two channels into which the wavelength-shifting fibers are inserted [4].
The fibers are read out by mm2 silicon photomultipliers located at each end, except in special cases where only one side is utilized.222The SiPMs are Hamamatsu S13360-2050VE devices with a 50 m pixel size, the large size being chosen as a compromise between dynamic range and quantum efficiency. For readout simplification, two bars are glued together side-by-side to form a di-counter (Fig. 2). An acetal fiber guide bar (FGB) is attached to each end of a di-counter, into which the fibers are glued with their ends polished. (The fibers are not glued into the scintillator channels.) Each SiPM is mounted on a printed circuit board by the manufacturer [5] (called SiPM carrier boards, see Fig. 2)). The SiPMs are housed in small rectangular wells in a black anodized aluminum fixture called the SiPM mounting block, which is precisely registered to the FGB by a pair of sleeves. A small circuit board, the counter motherboard (CMB), has spring-loaded “pogo” pins that gently press the SiPMs against the fibers and provide electrical contact to the SiPM. It provides the SiPM bias voltage and sends the four SiPM output signals of each di-counter end to a readout board via an HDMI cable. The CMB also has two flasher LEDs, one for each bar, and a single thermometer centered on the board.333The thermometer is a UMW DS18B20U digital temperature sensor with a range from -55∘C to 125∘C and an accuracy of C. This thermometer’s temperature measurement is used as a proxy for the temperature of the four SiPMs serviced by the CMB.
@@ 3 Readout of the Cosmic Ray Veto
The FEBs receive power through CAT6 cables from a readout controller (ROC), which also furnishes the clock timing information. In the normal readout of the CRV, upon receiving a trigger, each ROC accumulates data from up to 24 FEBs and sends it to the Data Transfer Controllers (DTCs), which serve as the interface between the data acquisition hardware and the online computer farm.
@@
Each FEB has eight Analog Front End (AFE) chips that serve to amplify, shape, and digitize (12-bit ADC) the signals from eight SiPMs at 80 MSPS (Mega samples per second or every 12.5 ns)[7]. The digitized signals from the eight AFEs are sent to four FPGAs (each FPGA handling 16 channels). The FPGAs perform a set of functions — such as zero-suppression — that were not used in this work. The data from the FPGAs are sent to a buffer memory where they await a trigger to be read out.
@@
The data acquisition process at the Wideband hall consisted of two phases: a two-minute live spill period during which triggers are accepted by the FEB, followed by a one-minute off-spill period dedicated to transmitting the recorded data to the PC for storage. Before the onset of each spill, slow control information was retrieved from the FEBs and written to the raw data file. This information included the settings of the AFEs, bias settings for each SiPM, CMB and FEB temperature, and the AFE gain settings.444The FEBs have an on-board temperature sensor.
During the live spill, the digitized SiPM-generated data were continuously written to the buffer memory on the FEB (at 80 MSPS). Upon receiving a trigger, data before and after the trigger time were transmitted. Typically 750 ns of data were read out before the trigger (pre-signal region — see Fig 4), allowing the pedestal to be determined as well as an occasional dark count from the SiPM noise, the latter being used in the calibration process. A total of 127 ADC samples were recorded for each of the 64 FEB channels. This sequence of data initiated by a single trigger is called an event. The event rate was approximately 3 Hz for the cosmic-ray trigger described below. For the LED flasher trigger, also described below, the event rate was 1 kHz.
@@ 4.1 Calibration Overview at a Stable Temperature
In the pre-signal region, the data in each channel undergoes a scan for dark pulses. An example dark pulse can be seen in Fig 4. The pulses undergo pedestal subtraction and are fitted using a Gumbel function given by the equation:
where is the time of the pulse peak, is the pulse height, and is the pulse area. Note that the Gumbel function was selected empirically based on the quality of pulse fitting.
The fitting process determines the dark count pulse areas, the distribution of which can be seen in Fig. 5, showing distinct peaks corresponding to the pulse areas of 1 PE, 2 PEs, and so forth.555Note that strictly speaking SiPMs produce a signal that is proportional to the number of hit pixels, not PEs. At low light levels the correspondence between the two is quite good. In this paper, we will use PEs rather than hit pixels. The single photoelectron peak (SPE) in the distribution of Fig. 5 is used to produce the SPE calibration constant that translates the SiPM waveform area into the reconstructed number of photoelectrons.666Due to a low cross-talk rate and hence limited statistics for the photoelectron peaks greater than 1 PE, only the SPE peak is utilized to establish the SPE calibration constant. The stability of the SPE calibration constant across a broader range of PEs is verified through LED flasher runs, as shown in Fig. 6, where Reco PEs stands for peaks in the SiPM output as fitted by the Gumbel function.777When determining pedestals we have subtracted 0.13 ADC counts from their mean values in order to get a more perfect linear fit between the reconstructed PEs and the expected PEs.
The calibration procedure described above works well for reasonably stable temperature environments and has been used in test beam experiments to study the CRV performance [8]. However, when efforts to precisely study the CRV aging properties over several years were attempted, it became clear that the temperature dependence of the CRV performance could not be ignored.888With the calibration scheme described here, precise aging measurements of the CRV counters are possible in the non-temperature controlled environment in which the counters are stored; these will be reported in a forthcoming publication. Below we describe the updated calibration procedure for equalizing the CRV response to a temperature that was primarily derived to make aging measurements of the CRV counters, but will also be valuable in the operation of the Mu2e experiment.
@@ 4.2 Calibrating for the Temperature Dependence
The SPE calibration constants and photo-detection efficiencies are impacted by the temperatures of both the SiPMs (as measured by the thermometer on the CMBs) and FEB ( and ), the former because the SiPMs have temperature dependent breakdown voltage, and the latter because of the bias provided by the FEBs to the SiPMs is slightly temperature dependent and therefore not exactly at its set value. In addition, the gain of the AFE amplifier/digitizers have a temperature dependence. Consequently, a temperature correction is essential to adjust the SPE constants and photo-detection efficiencies to reference temperatures, chosen to be: 20∘C for the SiPM and 40∘C for the FEB (roughly, their normal operating temperatures). This adjustment follows a specific procedure, as outlined in Fig. 7 and described in detail below.
The flow diagram of the calibration procedure is given in Fig. 7. In Step 1, described in Sec. 4.3, dark pulses are collected, fitted, and the pulse areas are extracted, . The SiPMs are not noisy, and hence the collection of sufficient numbers of dark pulses for a precise measurement of the SPE calibration constant might occur on a time scale in which the temperature is not constant. Hence, the dark pulses must first be corrected to the reference temperature before the SPE calibration constant, , is extracted from the single photoelectron peak. In Step 2, described in Sec. 5, the calibrations are applied to a measured signal pulse area, , at an arbitrary temperature, . must first be corrected to the proper value at the measurement temperature, . It can then be applied to the measured pulse area, , to estimate the number of photoelectrons observed. To compare the performance of the CRV over time, Step 3 of Fig. 7, also described in Sec. 5, can be applied to rescale the measured PE value to the reference temperature. In this way, CRV performance over time can be compared even if the temperature is not stable.
@@
Changes in over-voltage due to temperature fluctuations, in turn, lead to corrections of the SPE constants ( in Table 1). The SPE constants are also influenced by a gain change in the AFE ( in Table 1).999Parameters and depend on the AFE gain settings, and the values in the table are only valid for the specific experimental setup. Finally, changes to the over-voltage also require corrections to the light yield ( in Table 1).
@@ 4.3 Obtaining the Reference Calibration
Each pulse collected during the dark count scan undergoes an adjustment to its measured pulse area to its value at the reference temperatures and . For each pulse, the deviation of the over-voltage due to the temperature difference is calculated:
@@
These temperature-corrected pulse areas are then histogrammed in a calibration plot, and the temperature-corrected SPE constants, , are determined using the same approach as described in Sec. 4.1 and shown in Fig. 5. The parameters in Eq. 3 were obtained for temperature corrections of the SPE calibration constant, which is the pulse area of a single photoelectron. Therefore, this equation is only valid for pulse areas of single PEs, and only the single-PE peak of the temperature-corrected calibration plot is used to determine the SPE calibration constant.
@@ 5 Reconstructing the Light Yield
SiPM signal pulses are pedestal-subtracted and fitted with a Gumbel function in the same way as the dark counts shown in Eq. 1.101010Note that the temperature dependence of the pedestals were found to be negligible. Their pulse areas, peak times, leading-edge times, and width parameters are recorded.
@@
The second step is adjusting the calibration SPE constant from its value at the reference temperatures () to a calibration constant at the actual temperatures when the pulse was recorded (). This is done using the over-voltage change from Eq. 4 and the gain change due to the FEB temperature is an additional adjustment (analogous to Eq. 3 above but adjusting the calibration constant instead of the pulse area):
@@ 6 Determination of Temperature Correction Parameters
To determine the temperature corrections, LED data were collected and studied at Fermilab’s SiDet facility using the environmental chamber illustrated in Fig. 9. We measured the temperature dependence of the SiPMs, with the bias provided by a FEB situated outside of the chamber. During the studies described in this section, the AFE chip gains were uniformly set to their default operating value.111111The AFE gain was set to 33 dB, for a voltage gain of 45. The SPE-related parameters would change should the AFE chip gain vary.
@@ 6.1 SiPM Temperature Corrections
The configuration illustrated in Fig. 10 was used to measure the temperature dependence of the SiPMs. Bare SiPMs (uncoupled to fibers) were placed inside the environmental chamber, each group of four serviced by one CMB. The average of the CMB sensor readings was used as a proxy for the SiPM temperatures, and was found to agree well with the set temperature of the oven. An outside CMB flashed LED light through a transparent window. Two FEBs were employed in making the measurements. One was used to provide the SiPM biases and read out the SiPM signals; the other was used to flash the LED mounted on the outside CMB. Both were exposed to a relatively constant room temperature. The SiPMs’ dark pulses and their responses to the LED flashes were recorded for the chamber temperature range of -10∘C to 40∘C. For each temperature setting, the SiPM response was taken at a biases between 2 V and 6 V over-voltage (as calculated from the datasheet values provided by Hamamatsu). The applied bias voltages were recorded as .
@@ 6.1.1 SiPM Breakdown Voltage and Temperature Dependence
Bias scans were performed at each temperature setting of the environmental chamber, and the SPE response was recorded. The results, given in Fig. 11, show a linear dependence on the bias. The breakdown voltage is determined by extrapolating the linear fits to the point where SPE = 0. The extracted breakdown voltages for all 58 SiPMs and each temperature are shown in the top plot of Fig. 12.121212Fluctuations in the breakdown voltages of SiPMs 30–40 are attributed to production effects. Note that these same variations were observed in independent quality control measurements of the SiPMs. The SiPMs used in these studies were selected randomly from several batches received from the manufacturer and some variation is not surprising. Note that although 64 SiPMs were inside the chamber, 6 of them has issues that prevented their proper readout and were not used in the analysis.
The well-known increase in breakdown voltage with increasing temperature is apparent from Fig. 12. Averaging over all SiPMs, the extracted breakdown voltage increases at the rate of mV/K (Fig. 13 right). This is parameter in Table 1. It is worth noting that this value agrees with the specifications of 54 mV/K provided by Hamamatsu [5].
@@ 6.1.2 SiPM Single Photoelectron Yield Dependence on Over-Voltage
The temperature dependence of the breakdown voltage affects the photoelectron yield of the SiPMs, and hence the SPE calibration constants. That dependence was measured to be , from the fit parameters given in Fig. 11). This is parameter in Table 1. It is specific to the AFE chip gain setting we use for the CRV. The rate exhibits variations due to SiPM and AFE chip gain differences, but notably, the rate of SPE increase remains independent of the operating temperature over the range tested (Fig. 14). Since the SiPM-to-SiPM variation is small, using the average value of yields satisfactory results, as will be shown in Sec. 6.3 and Sec. 7.

### v2

@@ 1 Introduction
The Mu2e experiment intends to search for the conversion of a negatively charged muon into an electron in the presence of an aluminum nucleus: [1]. In the Standard Model, this process is forbidden. In the extended Standard Model with non-zero neutrino masses, the ratio of conversion to capture is at an experimentally unobservable value of [2]. Hence, any observation of this process would be unambiguous evidence of new physics.
@@ 2 The Cosmic Ray Veto Detector Design
The CRV scintillator bars are grouped in modules, each module consisting of 64 bars, 16 per layer, with aluminum sheets of 3/8” (9.525 mm) thickness placed between each scintillator layer. The entire CRV consists of 83 total modules. Each scintillator bar is mm2 in cross section, with lengths ranging from 0.985 m to 6.900 m.111The bars have a polystyrene base doped with 1% PPO, and 0.03% POPOP. The bars have a thin TiO2 reflective coating and two channels into which the wavelength-shifting fibers are inserted 222Kuraray America, Inc.. The fibers have diameters of 1.4 mm and 1.8 mm.
The fibers are read out by mm2 silicon photomultipliers located at each end, except in special cases where only one side is utilized. 333The SiPMs are Hamamatsu S13360-2050VE devices with a 50 m pixel size, the large size being chosen as a compromise between dynamic range and quantum efficiency. For readout simplification, two bars are glued together side-by-side to form a di-counter (Fig. 2). An acetal fiber guide bar (FGB) is attached to each end of a di-counter, into which the fibers are glued with their ends polished. (The fibers are not glued into the scintillator channels.) Each SiPM is mounted on a printed circuit board by the manufacturer 444Hamamatusu Corp. (called SiPM carrier boards, see Fig. 2). The SiPMs are housed in small rectangular wells in a black anodized aluminum fixture called the SiPM mounting block, which is precisely registered to the FGB by a pair of sleeves. A small circuit board, the counter motherboard (CMB), has spring-loaded “pogo” pins that gently press the SiPMs against the fibers and provide electrical contact to the SiPM. It provides the SiPM bias voltage and sends the four SiPM output signals of each di-counter end to a readout board via an HDMI cable. The CMB also has two flasher LEDs, one for each bar, and a single thermometer centered on the board.555The thermometer is a UMW DS18B20U digital temperature sensor with a range from -55∘C to 125∘C and an accuracy of C. This thermometer’s temperature measurement is used as a proxy for the temperature of the four SiPMs serviced by the CMB.
@@ 3 Readout of the Cosmic Ray Veto
The FEBs receive power through CAT6 cables from a readout controller (ROC), which also furnishes the clock timing information. In the normal readout of the CRV, upon receiving a trigger, each ROC accumulates data from up to 24 FEBs and sends it, by means of optical links, to the Data Transfer Controllers (DTCs), which serve as the interface between the data acquisition hardware and the online computer farm.
@@
Each FEB has eight Analog Front End (AFE) chips666AFE5807, Texas Instruments, Inc., https://www.ti.com/product/AFE5807 that serve to amplify, shape, and digitize (12-bit ADC) the signals from eight SiPMs at 80 MSPS (Mega samples per second or every 12.5 ns). The digitized signals from the eight AFEs are sent to four FPGAs (each FPGA handling 16 channels). The FPGAs perform a set of functions — such as zero-suppression — that were not used in this work. The data from the FPGAs are sent to a buffer memory where they await a trigger to be read out.
@@
The data acquisition process at the Wideband hall consisted of two phases: a two-minute live spill period during which triggers are accepted by the FEB, followed by a one-minute off-spill period dedicated to transmitting the recorded data to the PC for storage. Before the onset of each spill, slow control information was retrieved from the FEBs and written to the raw data file. This information included the settings of the AFEs, bias settings for each SiPM, CMB and FEB temperature, and the AFE gain settings.
During the live spill, the digitized SiPM-generated data were continuously written to the buffer memory on the FEB (at 80 MSPS). Upon receiving a trigger, data before and after the trigger time were transmitted. Typically 750 ns of data were read out before the trigger (pre-signal region — see Fig. 4), allowing the pedestal to be determined, and capturing an occasional dark count from the SiPM noise, the latter being used in the calibration process. A total of 127 ADC samples were recorded for each of the 64 FEB channels. This sequence of data initiated by a single trigger is called an event. The event rate was approximately 3 Hz for the cosmic-ray trigger described below. For the LED flasher trigger, also described below, the event rate was 1 kHz.
@@ 4.1 Calibration Overview at a Stable Temperature
In the pre-signal region, the data in each channel undergoes a scan for dark pulses. An example dark pulse can be seen in Fig. 4. The pulses undergo pedestal subtraction and are fitted using a Gumbel function given by the equation:
where is the time of the pulse peak, is the pulse height ( is Euler’s number), and is the pulse area. Note that the Gumbel function [5] was selected empirically based on the quality of pulse fitting.
The fitting process determines the dark count pulse areas, the distribution of which can be seen in Fig. 5, showing distinct peaks corresponding to the pulse areas of 1 PE, 2 PEs, and so forth.777Note that strictly speaking SiPMs produce a signal that is proportional to the number of hit pixels, not PEs. At low light levels the correspondence between the two is quite good. In this paper, we will use PEs rather than hit pixels. The single photoelectron peak (SPE) in the distribution of Fig. 5 is used to produce the SPE calibration constant that translates the SiPM waveform area into the reconstructed number of photoelectrons.888Due to a low cross-talk rate and hence limited statistics for the photoelectron peaks greater than 1 PE, only the SPE peak is utilized to establish the SPE calibration constant. The stability of the SPE calibration constant across a broader range of PEs is verified through LED flasher runs, where the LED flashers can be tuned in such a way that they produce only low PE pulses. For these kinds of signals, individual PE peaks can be easily separated as shown in Fig. 6, where Reco PEs stands for peaks in the SiPM output as fitted by the Gumbel function.999When determining pedestals we have subtracted 0.13 ADC counts from their mean values in order to get a more perfect linear fit between the reconstructed PEs and the expected PEs.
The calibration procedure described above works well for reasonably stable temperature environments and has been used in test beam experiments to study the CRV performance [6]. However, when efforts to precisely study the CRV aging properties over several years were attempted, it became clear that the temperature dependence of the CRV performance could not be ignored.101010With the calibration scheme described here, precise aging measurements of the CRV counters are possible in the non-temperature controlled environment in which the counters are stored; these will be reported in a forthcoming publication. Below we describe the updated calibration procedure for equalizing the CRV response to a temperature that was primarily derived to make aging measurements of the CRV counters, but will also be valuable in the operation of the Mu2e experiment.
@@ 4.2 Calibrating for the Temperature Dependence
The SPE calibration constants and photo-detection efficiencies are impacted by the temperatures of both the SiPMs (as measured by the thermometer on the CMBs) and FEB ( and ), the former because the SiPMs have temperature dependent breakdown voltage, and the latter because of the bias provided by the FEBs to the SiPMs is slightly temperature dependent and therefore not exactly at its set value. In addition, the gain of the AFE amplifier/digitizers have a temperature dependence. Consequently, a temperature correction is essential to adjust the SPE constants and photo-detection efficiencies to reference temperatures, chosen to be: 20∘C for the SiPM and 40∘C for the FEB (roughly, their usual operating temperatures). This adjustment follows a specific procedure, as outlined in Fig. 7 and described in detail below.
In Step 1, described in Sec. 4.3, dark pulses are collected, fitted, and the pulse areas are extracted, . The SiPMs are not noisy, and hence the collection of sufficient numbers of dark pulses for a precise measurement of the SPE calibration constant might occur on a time scale in which the temperature is not constant. Hence, the dark pulses must first be corrected to the reference temperature before the SPE calibration constant, , is extracted from the single photoelectron peak. In Step 2, described in Sec. 5, the calibrations are applied to a measured signal pulse area, , at an arbitrary temperature, . must first be corrected to the proper value at the measurement temperature, . It can then be applied to the measured pulse area, , to estimate the number of photoelectrons observed. To compare the performance of the CRV over time, Step 3 of Fig. 7, also described in Sec. 5, can be applied to rescale the measured PE value to the reference temperature. In this way, CRV performance over time can be compared even if the temperature is not stable.
@@
Changes in over-voltage due to temperature fluctuations, in turn, lead to corrections of the SPE constants ( in Table 1). The SPE constants are also influenced by a gain change in the AFE ( in Table 1).111111Parameters and depend on the AFE gain settings, and the values in the table are only valid for the specific experimental setup. Finally, changes to the over-voltage also require corrections to the light yield ( in Table 1).
@@ 4.3 Obtaining the Reference Calibration
Each pulse collected during the dark count scan undergoes an adjustment of its measured pulse area to its value at the reference temperatures and . For each pulse, the deviation of the over-voltage due to the temperature difference is calculated:
@@
These temperature-corrected pulse areas are then histogrammed in a calibration plot, and the temperature-corrected SPE constants, , are determined using the same approach as described in Sec. 4.1 and shown in Fig. 5. The parameters in Eq. 3 were obtained for temperature corrections of the SPE calibration constant, which is the pulse area of a single photoelectron. Therefore, this equation is only valid for pulse areas of single PEs, and only the first PE peak of the temperature-corrected calibration plot is used to determine the SPE calibration constant.
@@ 5 Reconstructing the Light Yield
SiPM signal pulses are pedestal-subtracted and fitted with a Gumbel function, shown in Eq. 1, in the same way as the dark counts.121212Note that the temperature dependence of the pedestals were found to be negligible. Their pulse areas, peak times, leading-edge times, and width parameters are recorded.
@@
The second step is adjusting the calibration SPE constant from its value at the reference temperatures () to a calibration constant at the actual temperatures when the pulse was recorded (). This is done using the over-voltage change from Eq. 4 and the gain change due to the FEB temperature as an additional adjustment (analogous to Eq. 3 above but adjusting the calibration constant instead of the pulse area):
@@ 6 Determination of Temperature Correction Parameters
To determine the temperature corrections, LED data were collected and studied at Fermilab’s SiDet facility using the environmental chamber illustrated in Fig. 9. We measured the temperature dependence of the SiPMs, with the bias provided by a FEB situated outside of the chamber. During the studies described in this section, the AFE chip gains were uniformly set to their default operating value.131313The AFE gain was set to 33 dB, for a voltage gain of 45. The SPE-related parameters would change should the AFE chip gain vary.
@@ 6.1 SiPM Temperature Corrections
The configuration illustrated in Fig. 10 was used to measure the temperature dependence of the SiPMs. Bare SiPMs (uncoupled to fibers) were placed inside the environmental chamber, each group of four serviced by one CMB. The average of the CMB sensor readings was used as a proxy for the SiPM temperatures, and was found to agree well with the set temperature of the chamber. An outside CMB flashed LED light through a transparent window. Two FEBs were employed in making the measurements. One was used to provide the SiPM biases and read out the SiPM signals; the other was used to flash the LED mounted on the outside CMB. Both were exposed to a relatively constant room temperature. The SiPMs’ dark pulses and their responses to the LED flashes were recorded for the chamber temperature range from -10∘C to 40∘C. For each temperature setting, the SiPM response was taken at a biases between 2 V and 6 V over-voltage (as calculated from the datasheet values provided by Hamamatsu). The applied bias voltages were recorded as .
@@ 6.1.1 SiPM Breakdown Voltage and Temperature Dependence
Bias scans were performed at each temperature setting of the environmental chamber, and the SPE response was recorded. The results, given in Fig. 11, show a linear dependence on the bias. The breakdown voltage is determined by extrapolating the linear fits to the point where SPE = 0. The extracted breakdown voltages for all 58 SiPMs and each temperature are shown in the top plot of Fig. 12.141414Fluctuations in the breakdown voltages of SiPMs 30–40 are attributed to production effects. Note that these same variations were observed in independent quality control measurements of the SiPMs. The SiPMs used in these studies were selected randomly from several batches received from the manufacturer and some variation is not surprising. Note that although 64 SiPMs were inside the chamber, 6 of them has issues that prevented their proper readout and were not used in the analysis. The four SiPMs shown as examples in Fig. 11 were selected arbitrarily; the same devices are used in subsequent figures to illustrate example results.
The well-known increase in breakdown voltage with increasing temperature is apparent from Fig. 12. Averaging over all SiPMs, the extracted breakdown voltage increases at the rate of mV/K (Fig. 13 right). This is parameter in Table 1. It is worth noting that this value agrees with the specifications of 54 mV/K provided by Hamamatsu.
@@ 6.1.2 SiPM Single Photoelectron Yield Dependence on Over-Voltage
The temperature dependence of the breakdown voltage affects the photoelectron yield of the SiPMs, and hence the SPE calibration constants. That dependence was measured to be , from the fit parameters given in Fig. 11. This is parameter in Table 1. It is specific to the AFE chip gain setting we use for the CRV. The rate exhibits variations due to SiPM and AFE chip gain differences, but notably, the rate of SPE increase remains independent of the operating temperature over the temperature range tested (Fig. 14). Since the SiPM-to-SiPM variation is small, using the average value of yields satisfactory results, as will be shown in Sec. 6.3 and Sec. 7.